OpenStudy (anonymous):

What is the slope of the line that passes through the pair of points for (-5/3, -1), (-2, 9/2) please

OpenStudy (anonymous):

\[Slope = \frac{ (y _{2}-y _{1}) }{(x _{2}-x _{1}) }\] y2= 9/2 y1= -1 x2= -2 x1= -5/3 Substitute the values and simplify. Watch your signs.

OpenStudy (anonymous):

Hello Ryan, Thank you. I think that is my problem the signs keep getting me mixed up

OpenStudy (anonymous):

Yeah, it's easy to keep mixed up with signs when finding the slope. Did you get the answer?

OpenStudy (anonymous):

I am trying to work it out now.

OpenStudy (anonymous):

y= 5.5 I think

OpenStudy (anonymous):

x= -1/3

OpenStudy (anonymous):

Am I even close to the answer

OpenStudy (anonymous):

Yes, thats correct! \[\frac{ 5.5 }{ -1/3 }\] Good job

OpenStudy (anonymous):

Thank you so much. smiley faces

OpenStudy (anonymous):

Have a great day

OpenStudy (anonymous):

Yep, you too, glad I could help